OpenStudy (anonymous):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept.

OpenStudy (anonymous):

Using the slope-intercept form formula you can find right away the coords of the y-intercept; \[y=mx+b\] Where m represents the slop and b represents the y-intercept. \[y=-2x-3\] The y-intercept would be the value of y when x=0, and similarly the x-intercept would be the value of x when y=0. So, Let's do y-intercept first; \[y=-2(0)-3\] \[y=0-3\]\[y=-3\] So when x is 0, y is -3. Using;\[(x, y)\] The y-intercept is; \[(0, -3)\] Now to find the x-intercept; \[0=-2x-3\]\[0+3=-2x-3+3\]\[3=-2x\]\[\frac{ 3 }{ -2 }=\frac{ -2x }{ -2 }\]\[-1.5=x\] So, when y = 0, x = -1.5. Using;\[(x, y)\]The x-intercept would be; \[(-1.5, 0)\]